### v1

@@ Abstract
We study natural privacy filters, which enable the exact composition of differentially private (DP) mechanisms with adaptively chosen privacy characteristics. Earlier privacy filters consider only simple privacy parameters such as Rényi-DP or Gaussian DP parameters. Natural filters account for the entire privacy profile of every query, promising greater utility for a given privacy budget. We show that, contrary to other forms of DP, natural privacy filters are not free in general. Indeed, we show that only families of privacy mechanisms that are well-ordered when composed admit free natural privacy filters.
@@ 1 Introduction
Differential Privacy (DP) [dwork2006calibrating] is a rigorous notion of data privacy, used to protect against a host of different privacy attacks, including membership inference [homer2008resolving, dwork2015robust, carlini2022membership] and data reconstruction [dinur2003revealing, dwork2007price, carlini2021extracting, balle2022reconstructing, dick2023confidence, nasr2023scalable]. As such, there is a sustained push to integrate and deploy DP in a large number of applications [opendp-registry], from machine learning (ML) and artificial intelligence (AI) models [kairouz2021practical, sinha2025vaultgemma], to analytics collection [rogers2020linkedin, apple-emoji, ding2017collecting], to data releases [adeleye2023publishing, abowd20222020], as well as in general libraries and systems supporting those applications [gaboardi2020programming, wilson2020differentially, jax-privacy2022github, tholoniat2024cookie, ghazi2025differential]. This practical DP push also motivates important theoretical progress.
A prominent direction of theoretical improvement comes from alternative definitions of DP, that enable tighter privacy loss accounting under composition and improve the privacy-accuracy tradeoff. The most successful definitions include $(\varepsilon,\delta)$-DP [dwork2006our] (known as approximate DP), the Rényi divergence family of DP definitions such as zCDP [dwork2016concentrated, bun2016concentrated] and RDP [mironov2017renyi], and privacy definitions that generalize approximate DP by considering all $(\varepsilon,\delta)$-DP values compatible with a given mechanism, such as the privacy loss distribution (PLD) [sommer2019privacy] and $f$-DP [dong2022gaussian]. The latter definitions, PLD and $f$-DP, are theoretically appealing as they tightly capture resilience to membership inference attacks framed as hypothesis tests [3, dong2022gaussian]. However, composition under these definitions is more complex and can be computationally difficult, with analyses of specific algorithms that rely on numerical approaches to handle composition in practice [koskela2020computing, gopi2021numerical].
@@
Two other theoretical improvements enable new forms of composition that are necessary requirements for practical, real world DP usage. First, concurrent composition [vadhan2021concurrent] supports composition of DP mechanisms running over multiple rounds of interaction with the user—such as the sparse vector technique [dwork2009complexity]—that are running in parallel. This is required in practical systems which typically have to handle queries from multiple parties simultaneously, and where it is not possible to enforce a sequential ordering of the queries. Even individual users may want to run DP computations concurrently with long-running, multi-step mechanism (e.g., [kostopoulou2023turbo]).
Second, adaptive composition, the main topic of this paper, enables the composition of DP mechanisms with adaptively chosen privacy parameters (as opposed to regular composition, in which these parameters have to be known in advance). Adaptive composition is formalized through the notion of privacy filters [RRUV16], which accept DP mechanisms as long as their composition remains within a bound fixed in advance (more details in §2). Such adaptive composition is necessary to support long running applications, in which analysts can adaptively query a system with mechanisms and DP parameters of their choice [lecuyer2019privacy, kostopoulou2023turbo, kuchler2024cohere]. Furthermore, filters enable crucial privacy accounting savings through advanced versions of parallel composition [lecuyer2019privacy], as well as individual (personalized) DP [ebadi2015differential], which enables data dependent privacy accounting in ML model training [feldman2021individual, yu2022individual] and measurement systems [tholoniat2024cookie, ghazi2025differential].
A key question in DP composition is whether these various improvements combine with each other. Ideally these more flexible forms of composition come for free: that is, that can we apply composition results under concurrent mechanisms or adaptive budgets using privacy filters, without a penalty in the final privacy loss after composition. Flexible composition for free is important not only because it is efficient privacy-wise, but also because it enables reusing all existing composition results, including tight analyses for specific mechanisms and empirical computational results.
As it turns out, concurrent composition is indeed free for all DP definitions [vadhan2023concurrent, lyu2022composition]. Moreover, concurrent and adaptive filter composition reduces to adaptive filters [haney2023concurrent], and thus has the same composition cost. Therefore, the full expressivity of both concurrency and adaptivity hinges on free privacy filter composition for different notions of DP. Filters, however, are not as well understood. We know that adaptive composition under privacy filters is free for $\varepsilon$-DP [RRUV16] and Rényi-based DP definitions [FZ21, lecuyer2021practical] (zCDP, RDP). Filters are also free for Gaussian mechanisms under $f$-DP and PLD accounting [ST22, KTH22]. However, for $(\varepsilon,\delta)$-DP, the best known result for adaptive composition with filters is only asymptotically free, with worse constants than non-adaptive composition [WRRW23]. Even less is known about filters under $f$-DP and PLD. In particular, we consider natural filters, which leverage lossless composition of the privacy characteristics of individual queries. Natural filters can be realized via PLDs, privacy profiles, or $f$-DP accounting. In this paper, we make three important contributions towards fully understanding natural privacy filters:
@@
  • In §4 we apply this theory to the natural filter and, surprisingly, we show that adaptive composition is NOT free under the PLD nor $f$-DP notions of privacy (see Corollary˜1).
  • Finally, we show in §6 that filters are only free for very restricted families of mechanisms: those that are closed under composition, and admit totally ordered tradeoff functions.
Our work also includes some minor contributions including that PLDs can be assigned a natural ordering satisfying a completeness property enabling one to take a supremum of PLDs (see Proposition˜11). As far as we are aware this property has not been documented in the privacy accounting literature. We alo document in §5 some families of queries and budgets that do and do not admit free filters.
@@ 2 Preliminaries
Differential privacy is the study of randomized algorithms (called mechanisms) $M:\mathcal{D}\to\mathcal{Y}$ where $\mathcal{D}$ denotes some space of datasets containing sensitive personal information. Differential privacy requires that an adversary should not gain much statistical advantage for inferring whether the mechanism was run on dataset $D_{1}\in\mathcal{D}$ or a second dataset $D_{2}\in\mathcal{D}$ when the datasets are close to each other in some sense. That is, the distributions of $M(D_{1})$ and $M(D_{2})$ should be difficult to distinguish from one another in the hypothesis testing sense and this should hold uniformly for any close datasets. The measure of closeness for datasets is generally defined by a neighbouring relation but this turns out not to be relevant to our work. In fact, our conclusions hold for an arbitrary pair of datasets. For this reason, we fix throughout a single arbitrary pair of datasets $(D_{1},D_{2})\in\mathcal{D}\times\mathcal{D}$ and view the privacy characteristics of a mechanism $M$ in terms of the hypothesis testing problem $M(D_{1})$ vs. $M(D_{2})$.
@@
It is natural in general to view the privacy characteristics of any private mechanism as a hypothesis testing problem [3, DRS22]. As such it is desirable to furnish hypothesis testing problem $(P,Q)$ with some kind of ordering so that we can compare the privacy guarantees offered by various mechanisms. We also require an ordering so that we can make sense of privacy filters. Fortunately testing pairs can be equipped with a very natural order, namely the Blackwell order.
@@
For a pair of distributions $(P,Q)$ defined on a common probability space $\Omega$ and a second pair $(P^{\prime},Q^{\prime})$ defined on $\Omega^{\prime}$, we say that $(P^{\prime},Q^{\prime})$ dominates $(P,Q)$ in the Blackwell order, written $(P,Q)\preceq(P^{\prime},Q^{\prime})$, if we can find a Markov kernel $\phi$ from $\Omega^{\prime}$ to $\Omega$ such that $P=\phi P^{\prime}$ and $Q=\phi Q^{\prime}$. We write $(P,Q)\prec(P^{\prime},Q^{\prime})$ if $(P,Q)\preceq(P^{\prime},Q^{\prime})$ but not vice-versa.
Informally, the Blackwell order says that more information is available for distinguishing $P^{\prime}$ and $Q^{\prime}$ compared to $P$ vs $Q$. For mechanisms, we can interpret this as $(M(D_{1}),M(D_{2}))\preceq(M^{\prime}(D_{1}),M^{\prime}(D_{2}))$ when $M$ can be reconstructed from $M^{\prime}$ by introducing additional randomness through the kernel $\phi$. Note that the Blackwell order is a partial order.
Instead of a carrying around a pair of distributions to represent the privacy characteristics of a mechanism, we can consolidate them into the privacy loss distribution (PLD) [2]. As we will see the PLD possesses very desirable composition properties and is a practical basis for privacy loss computations. To formalize the PLD, we require likelihood ratios. Recall that, for a distribution $P$ absolutely continuous with respect to a distribution $Q$ ($P\ll Q$) on a probability domain $\Omega$, their likelihood ratio is the Radon-Nikodym derivative $\ell(\omega):=\frac{dP}{dQ}(\omega)$. However, in some cases, we would like to consider private mechanisms that violate absolute continuity and still construct the PLD. This can be achieved by taking any measure $\nu$ on $\Omega$ such that $P,Q\ll\nu$ (e.g. $\nu:=P+Q$) and setting $\ell(\omega):=\frac{dP}{d\nu}(\omega)/\frac{dQ}{d\nu}(\omega)$ ($\cdot/0:=\infty$). This construction of the likelihood ratio is unique only outside of a null-measure set [^1] More carefully, assume $P,Q\ll\nu,\nu^{\prime}$ and set $\mu:=\nu+\nu^{\prime}$ so that $P,Q\ll\nu,\nu^{\prime}\ll\mu$. By Radon–Nikodym and the chain rule for measures, we have, for $\omega\sim P$, that $\frac{dP}{d\nu}(\omega)/\frac{dQ}{d\nu}(\omega)=\frac{dP}{d\mu}(\omega)/\frac{dQ}{d\mu}(\omega)=\frac{dP}{d\nu^{\prime}}(\omega)/\frac{dQ}{d\nu^{\prime}}(\omega)$ with probability $1$., which is enough to construct a unique PLD. For this reason, we denote by $\frac{dP}{dQ}$ the likelihood ratio of $P$ with respect to $Q$ even when $P\not\ll Q$.
@@
Let $P$ and $Q$ be distributions on $\Omega$. The PLD of $(P,Q)$, denoted $\operatorname{PLD}(P\parallel Q)$, is the distribution over $\mathbb{R}\cup\{\infty\}$ of $\log(\frac{dP}{dQ}(\omega)),\omega\sim P$.
@@
A distribution $L$ on $\mathbb{R}\cup\{\infty\}$ is the PLD of some $(P,Q$) if and only if it satisfies $\mathop{\mathbb{E}}_{Z\sim L}[e^{-Z}]\leq 1$. In this case, one such pair is $(L,L^{\prime})$ where $L^{\prime}$ is the Esscher tilt of $L$, namely $dL^{\prime}(z):=e^{-z}dL(z)$ for $z\in\mathbb{R}$ and $L^{\prime}(\{-\infty\}):=1-\mathop{\mathbb{E}}_{Z\sim L}[e^{-Z}]$.
PLDs also inherit the Blackwell order: For PLDs $L$ and $L^{\prime}$, we say that $L\preceq L^{\prime}$ if there exist pairs $(P,Q)$ and $(P^{\prime},Q^{\prime})$ such that $L:=\operatorname{PLD}(P\parallel Q)$, $L^{\prime}=\operatorname{PLD}(P^{\prime}\parallel Q^{\prime})$, and $(P,Q)\preceq(P^{\prime},Q^{\prime})$. It is straightforward to show that this order is well-defined regardless of the underlying representations $(P,Q)$ and $(P^{\prime},Q^{\prime})$.
PLDs also form a monoid under convolution that is monotone in the Blackwell order.
@@
    The identically zero distribution Id is a PLD such that $L_{1}\oplus\text{Id}=\text{Id}\oplus L_{1}=L_{1}$;
@@
    $(L_{1}\oplus L_{2})\oplus L_{3}=L_{1}\oplus(L_{2}\oplus L_{3})$; and
@@
    If $L_{1}\preceq L^{\prime}_{1}$ and $L_{2}\preceq L^{\prime}_{2}$, then $L_{1}\oplus L_{2}\preceq L^{\prime}_{1}\oplus L^{\prime}_{2}$.
A highly prized property of PLDs is that the privacy characteristics of an adaptive mechanism corresponds to convolution of PLDs, provided that each component fixes a privacy budget in advance. This is important because the convolution of a long sequence of distributions can be computed very efficiently by applying the fast Fourier transform [koskela2020computing]. Note that we will require more than this to understand fully adaptive composition where the privacy bound of each component is itself adaptive.
@@
Let $M_{1}:\mathcal{D}\to\mathcal{Y}_{1}$ and $M_{2}:\mathcal{D}\times\mathcal{Y}_{1}\to\mathcal{Y}_{2}$ be adaptive mechanisms such that $\operatorname{PLD}(M_{1})\preceq L_{1}$ and $\operatorname{PLD}(M_{2}(\cdot;y_{1}))\preceq L_{2}$ for every $y_{1}\in\mathcal{Y}_{1}$. Then $\operatorname{PLD}(M_{1}\otimes M_{2})\preceq L_{1}\oplus L_{2}$. Moreover, if $\operatorname{PLD}(M_{1})\preceq\operatorname{PLD}(M^{\prime}_{1})$ and $\operatorname{PLD}(M_{2}(\cdot;y_{1}))\preceq\operatorname{PLD}(M^{\prime}_{2}(\cdot;y_{1}))$ for every $y_{1}\in\mathcal{Y}_{1}$, then $\operatorname{PLD}(M_{1}\otimes M_{2})\preceq\operatorname{PLD}(M^{\prime}_{1}\otimes M^{\prime}_{2})$.
Closely related to the PLD is the hockey-stick curve: a convex reparameterization of the privacy profile of a mechanism generalizing the classical notion of $(\varepsilon,\delta)$-DP.
@@
Again, for convenience, we write $H_{x}(M):=H_{x}(M(D_{1})\parallel M(D_{2}))$ and sometimes $h_{M}(x):=H_{x}(M)$ for the hockey-stick curve of a mechanism $M$. We will rely on hockey-stick curves in order to reason about adversarial behaviour of adaptive mechanisms. The hockey-stick curve is very closely related to the classical $(\varepsilon,\delta)$ form of differential privacy. In particular, by construction of the hockey-stick divergence a mechanism $M$ satisfies $(\varepsilon,\delta)$-DP exactly when $H_{e^{\varepsilon}}(M)\leq\delta$. In general, hockey-stick curves are characterized by a few simple conditions.
Proposition 4 ([4] Lemma 9).
A curve $h:\mathbb{R}^{\times}\to[0,1]$ is a hockey-stick curve of some pair $(P,Q)$ if and only if $h$ is convex and decreasing such that $\lim_{x\to 0}h(x)=1$ and $h(x)\geq 1-x$ for all $x\in\mathbb{R}^{\times}$.
Hockey-stick curves are also closely related to PLDs and can be fully recovered from the PLD via the following formula.

### v2

@@ Abstract
We study privacy filters, which enable privacy accounting for differentially private (DP) mechanisms with adaptively chosen privacy characteristics. We develop a general theory that characterizes the worst-case privacy loss of an interaction involving an analyst that respects some restrictions on what queries they may issue. We apply this theory to develop residue filters, which unifies existing privacy filters. We develop the Gaussian DP (GDP) residue filter, which strictly improves upon the naïve GDP filter. We also show that residue filters capture the natural filter, which promises greater utility by leveraging exact privacy accounting techniques. Earlier privacy filters consider only simple privacy parameters such as Rényi-DP or GDP parameters. Natural filters account for the entire privacy profile of every query, promising more efficient use of a given privacy budget. We show that, contrary to other forms of DP, natural privacy filters are not free in general. We present a characterization of when a family of private queries admits free natural filters for a given budget. In particular, only families of privacy mechanisms that are totally-ordered when composed admit free natural privacy filters with respect to an arbitrary privacy budget. Finally, we show that, while the natural approximate-DP filter can fail in the presence of adaptive adversary, it cannot fail too badly: the output remains approximate-DP with parameters at most poly-logarithmically worse than the intended privacy parameters.
@@ 1 Introduction
Differential Privacy (DP) Dwork et al. (2006b) is a rigorous notion of data privacy that protects against a host of attacks including membership inference Homer et al. (2008); Dwork et al. (2015); Carlini et al. (2022) and data reconstruction Dinur and Nissim (2003); Dwork et al. (2007); Carlini et al. (2021); Balle et al. (2022); Dick et al. (2023); Nasr et al. (2023). As such, there is a sustained push to integrate and deploy DP in a large number of applications OpenDP (2025), from machine learning (ML) and artificial intelligence (AI) models Kairouz et al. (2021); Sinha et al. (2025), to analytics collection Rogers et al. (2020); Apple (2017); Ding et al. (2017), to data releases Adeleye et al. (2023); Abowd et al. (2022), as well as in general libraries and systems supporting those applications Gaboardi et al. (2020); Wilson et al. (2020); Balle et al. (2025); Tholoniat et al. (2024); Ghazi et al. (2025). This practical DP push also motivates important theoretical progress.
A prominent direction of theoretical improvement comes from alternative quantifications of privacy, especially those that offer tighter or even lossless accounting of composition. The classical $(\varepsilon,\delta)$-DP Dwork et al. (2006a) (known as approximate DP) was first refined by forms of privacy derived from Rényi divergence including zCDP Dwork and Rothblum (2016); Bun and Steinke (2016) and RDP Mironov (2017). Later, exact forms of privacy accounting were developed including privacy profiles, privacy loss distributions (PLDs) Sommer et al. (2019), and $f$-DP Dong et al. (2019). These exact forms of accounting are practically and theoretically appealing as they tightly capture composition as well as resilience to membership inference attacks framed as hypothesis tests Wasserman and Zhou (2010); Dong et al. (2019). Exact composition, while more complex, can be computed efficiently to arbitrary precision by numerical methods Koskela et al. (2020); Gopi et al. (2021).
@@
Theoretical improvements enabling new forms of composition are also necessary for practical DP usage. Concurrent composition Vadhan and Wang (2021) supports composition of DP mechanisms running over multiple rounds of parallel interactions with the user—such as the sparse vector technique Dwork et al. (2009). This is required in systems that handle queries from multiple parties simultaneously where it is not possible to enforce a sequential ordering of the queries. Individual users may also run DP computations concurrently with a long-running, multi-step mechanism (see e.g. Kostopoulou et al. (2023)).
A natural question is whether developments in privacy accounting are compatible with new forms of composition. Ideally, more flexible forms of composition come for free, i.e. existing results for standard composition can be applied to flexibly composed mechanisms without additional privacy costs. Free flexible composition also enables reuse of existing results including tight analyses for specific mechanisms and empirical computational results. As it turns out, concurrent composition is indeed free for all common forms of privacy Vadhan and Zhang (2023); Lyu (2022). Specifically, Vadhan and Zhang Vadhan and Zhang (2023) show that for both $f$-DP (including PLD and approximate DP) and RDP (including zCDP) notions of privacy, concurrent composition can be seen as a post-processing of standard composition and thus has the same cost.
Another form of flexible composition and the subject of this work is fully adaptive composition, which concerns the composition of DP mechanisms with interactively chosen privacy costs. Standard composition on the other hand requires the privacy costs to be fixed in advance. This is insufficient for long-running applications in which analysts interactively re-assess the privacy budget allocated to queries in order to meet operational needs Lécuyer et al. (2019); Kostopoulou et al. (2023); Küchler et al. (2024). Fully adaptive analysis can also lead to substantial privacy loss savings through advanced versions of parallel composition Lécuyer et al. (2019) and individual (personalized) DP Ebadi et al. (2015), which enables data dependent privacy accounting in ML model training Feldman and Zrnic (2021); Yu et al. (2022) and measurement systems Tholoniat et al. (2024); Ghazi et al. (2025). Fully adaptive composition is formalized in terms of privacy filters Rogers et al. (2016). A privacy filter is an algorithm that monitors an interaction between an analyst and a data curator and permits queries as long as a privacy budget condition is maintained. The stopping condition should be as permissive as possible such that resulting interaction provably satisfies a meaningful privacy guarantee.
As with concurrent composition, a key research question is whether advanced privacy accounting techniques can be used to construct free privacy filters for fully adaptive composition. Haney et al. Haney et al. (2023) show that concurrent, adaptive composition reduces to adaptive composition. Therefore, the full expressivity of both concurrency and adaptivity hinges on the existence of free privacy filters for different notions of DP. However, privacy filters are not as well understood as concurrent composition. We know that adaptive composition under privacy filters is free for $\varepsilon$-DP Rogers et al. (2016) and Rényi-based DP Feldman and Zrnic (2021); Lécuyer (2021) (zCDP, RDP). The GDP filter, which applies exact ($f$-DP, PLD) accounting to the Gaussian mechanism is also free Smith and Thakurta (2022); Koskela et al. (2022). This filter was also recently extended to support approximate GDP budgets Tran et al. (2026). However, for $(\varepsilon,\delta)$-DP filters, the best known filter by Whitehouse et al. Whitehouse et al. (2023) extends the RDP filter from Feldman and Zrnic (2021) to approximate RDP and zCDP, enabling a filter based on lossy translation from $(\varepsilon,\delta)$-DP to approximate zCDP Bun and Steinke (2016). This yields an $(\varepsilon,\delta)$-DP privacy filter that is asymptotically free but with worse constants than non-adaptive composition due to translation loss. Even less is known about filters under exact forms of privacy. We study the natural filter in which the realized privacy costs of each query are composed by exact accounting and the realized total cost is compared to a fixed budget.
We make four important contributions towards fully understanding privacy filters:
@@
  • In §4, we show that existing privacy filters are subsumed by residue filters, which perform exact accounting under a residual update condition. We develop a new GDP residue filter that offers significant privacy savings compared to the standard GDP filter for some settings.
  • In §5 we apply our theory of constrained adversaries to calculate the exact privacy cost of an interaction with the natural filter. We show that the natural filter is universally free for any budget only for queries that are closed under composition and totally-ordered, in which case the natural filter is captured by a residue filter.
  • In §7, we consider the widely used natural $(\varepsilon,\delta)$-DP filter, i.e. the filter based on exact composition and $(\varepsilon,\delta)$-DP budget. We show that, while this filter can fail, it does not fail too badly: the interaction still satisfies $(\chi\cdot\varepsilon,\chi\cdot\delta)$-DP where $\chi$ is only poly-logarithmic in $1/\varepsilon,1/\delta$ and the number of rounds of adaptivity.
Our work also includes some minor contributions including that PLDs can be assigned a natural ordering satisfying a completeness property enabling one to take a supremum of PLDs (see Proposition˜11). As far as we are aware this property has not been documented in the privacy accounting literature. We also record in §6 a number of families of queries and budgets that do and do not admit free natural filters. Our key result showing the necessity of total-ordering for universally free natural filters is given in §8 and involves a privacy gap amplification technique that may be of general interest.
We note that a recent preprint independent and simultaneous to our work also studies natural filters Tran et al. (2026). That work states that a total-ordering condition is necessary and sufficient for free natural filters when the queries have unbounded support. Our work contains an analogous result for general queries. In particular, our result also applies to discretized finite-support queries that arise in practical privacy accounting. Our work also offers a fine-grained analysis of the cost of interacting with the natural filter for particular budgets.
@@ 2 Preliminaries
Differential privacy is the study of randomized algorithms (called mechanisms) $M:\mathcal{D}\to\mathcal{Y}$ where $\mathcal{D}$ denotes some space of datasets containing sensitive information. Differential privacy requires that an adversary should not gain much statistical advantage for inferring whether the mechanism was run on dataset $D_{1}\in\mathcal{D}$ or a second dataset $D_{2}\in\mathcal{D}$ when the datasets are close to each other in some sense. That is, the distributions of $M(D_{1})$ and $M(D_{2})$ should be difficult to distinguish from one another in the hypothesis testing sense and this should hold uniformly for any close datasets. The measure of closeness for datasets is determined by a neighbouring relation but this turns out not to be relevant to our work. In fact, our conclusions hold for an arbitrary pair of datasets. For this reason, we fix throughout a single arbitrary pair of datasets $(D_{1},D_{2})\in\mathcal{D}\times\mathcal{D}$ and view the privacy characteristics of a mechanism $M$ in terms of the hypothesis testing problem $M(D_{1})$ vs. $M(D_{2})$.
In general, the privacy characteristics of a mechanism may not be determined by a single pair. For the purposes of constructing filters, this can be handled by considering multiple dataset pairs that characterize some regime of the worst-case privacy characteristics of a mechanism. One can also smooth over the choices of datasets by considering dominating pairs, though this is lossy in general. See Zhu et al. (2022) for more discussion.
@@
It is natural in general to view the privacy characteristics of any private mechanism as a hypothesis testing problem Wasserman and Zhou (2010); Dong et al. (2019). As such it is desirable to furnish hypothesis testing problems $(P,Q)$ with some kind of ordering so that we can compare the privacy guarantees offered by various mechanisms. We also require an ordering so that we can make sense of natural privacy filters. Fortunately testing pairs can be equipped with a very natural order, namely the Blackwell order.
@@
For a pair of distributions $(P,Q)$ defined on a common probability space $\Omega$ and a second pair $(P^{\prime},Q^{\prime})$ defined on $\Omega^{\prime}$, we say that $(P^{\prime},Q^{\prime})$ dominates $(P,Q)$ in the Blackwell order, written $(P,Q)\preceq(P^{\prime},Q^{\prime})$, if we can find a Markov kernel $\kappa$ from $\Omega^{\prime}$ to $\Omega$ such that $P=\kappa P^{\prime}$ and $Q=\kappa Q^{\prime}$. We write $(P,Q)\prec(P^{\prime},Q^{\prime})$ if $(P,Q)\preceq(P^{\prime},Q^{\prime})$ but not vice-versa. If $(P,Q)\preceq(Q,P)\preceq(P,Q)$ the pair is called symmetric.
Informally, the Blackwell order says that more information is available for distinguishing $P^{\prime}$ and $Q^{\prime}$ compared to $P$ vs $Q$. For mechanisms, we can interpret this as $(M(D_{1}),M(D_{2}))\preceq(M^{\prime}(D_{1}),M^{\prime}(D_{2}))$ when $M$ can be reconstructed from $M^{\prime}$ by introducing additional randomness through the kernel $\kappa$. Note that the Blackwell order is a partial order.
Instead of a carrying around a pair of distributions to represent the privacy characteristics of a mechanism, we can consolidate them into a privacy loss distribution (PLD) Dwork and Rothblum (2016). As we will see, the PLD possesses very desirable composition properties and is a practical basis for privacy loss computations. To formalize the PLD, we require likelihood ratios. Recall that, for a distribution $P$ absolutely continuous with respect to a distribution $Q$ ($P\ll Q$) on a probability domain $\Omega$, their likelihood ratio is the Radon-Nikodym derivative $\frac{dP}{dQ}$. However, in some cases, we would like to consider private mechanisms that violate absolute continuity and still construct the PLD. In this case, we consider the Lebesgue decomposition of $P=P_{\ll}+P_{\bot}$ into absolutely continuous and singular components. For points in the support of the singular measure $P_{\bot}$, we assign likelihood $\infty$ and, for points outside, we may safely assign the likelihood $dP_{\ll}/dQ$. For simplicity, we denote this extended likelihood ratio as $dP/dQ$.
@@
Let $P$ and $Q$ be distributions on $\Omega$. The PLD of $(P,Q)$, denoted $\operatorname{PLD}(P\parallel Q)$, is the distribution over $\overline{\mathbb{R}}:=\mathbb{R}\cup\{\pm\infty\}$ of $\log(\frac{dP}{dQ}(\omega)),\omega\sim P$.
@@
A distribution $L$ on $\overline{\mathbb{R}}$ is the PLD of some $(P,Q$) if and only if it satisfies $\mathop{\mathbb{E}}_{Z\sim L}[e^{-Z}]\leq 1$. In this case, one such pair is $(L,L^{\prime})$ where $L^{\prime}$ is the Esscher tilt of $L$, namely $dL^{\prime}(z):=e^{-z}dL(z)$ for $z\in\mathbb{R}$ and $L^{\prime}(\{-\infty\}):=1-\mathop{\mathbb{E}}_{Z\sim L}[e^{-Z}]$.
PLDs also inherit the Blackwell order: For PLDs $L$ and $L^{\prime}$, we say that $L\preceq L^{\prime}$ if there exist pairs $(P,Q)$ and $(P^{\prime},Q^{\prime})$ such that $L:=\operatorname{PLD}(P\parallel Q)$, $L^{\prime}=\operatorname{PLD}(P^{\prime}\parallel Q^{\prime})$, and $(P,Q)\preceq(P^{\prime},Q^{\prime})$. It is straightforward to show that this order is well-defined regardless of the underlying representations $(P,Q)$ and $(P^{\prime},Q^{\prime})$. When equipped with the Blackwell order, PLDs also form a partially-ordered commutative monoid under convolution.
@@
    The identically zero distribution Id and the identically $\infty$ distribution, also denoted $\infty$, are PLDs such that $\text{Id}\preceq L_{1}\preceq\infty$, $L_{1}\oplus\text{Id}=L_{1}$, and $L_{1}\oplus\infty=\infty$;
@@
    $(L_{1}\oplus L_{2})\oplus L_{3}=L_{1}\oplus(L_{2}\oplus L_{3})$;
@@
    $L_{1}\oplus L_{2}=L_{2}\oplus L_{1}$; and
  5. 5.
    If $L_{2}\preceq L^{\prime}_{2}$, then $L_{1}\oplus L_{2}\preceq L_{1}\oplus L^{\prime}_{2}$.
A highly prized property of PLDs is that the privacy characteristics of an adaptive mechanism corresponds to convolution of PLDs, provided that each component fixes a privacy budget in advance. This is important because the convolution of a long sequence of distributions can be computed very efficiently by applying the fast Fourier transform Koskela et al. (2020). Note that we will require more than this to understand fully adaptive composition.
@@
Let $M_{1}:\mathcal{D}\to\mathcal{Y}_{1}$, $M^{\prime}_{1}:\mathcal{D}\to\mathcal{Y}^{\prime}_{1}$, $M_{2}:\mathcal{D}\times\mathcal{Y}_{1}\to\mathcal{Y}_{2}$, and $M^{\prime}_{2}:\mathcal{D}\times\mathcal{Y}^{\prime}_{1}\to\mathcal{Y}^{\prime}_{2}$ be adaptive mechanisms such that $\operatorname{PLD}(M_{1})\preceq\operatorname{PLD}(M^{\prime}_{1})$ and $\operatorname{PLD}(M_{2}(\cdot;y_{1}))\preceq\operatorname{PLD}(M^{\prime}_{2}(\cdot;y_{1}))$ for every $y_{1}\in\mathcal{Y}_{1}$. Then $\operatorname{PLD}(M_{1}\otimes M_{2})\preceq\operatorname{PLD}(M^{\prime}_{1}\otimes M^{\prime}_{2})$. Moreover, if $\operatorname{PLD}(M_{1})=L_{1}$ and $\operatorname{PLD}(M_{2}(\cdot;y_{1}))=L_{2}$ for every $y_{1}\in\mathcal{Y}_{1}$, then $\operatorname{PLD}(M_{1}\otimes M_{2})=L_{1}\oplus L_{2}$.
We also define a few important classes of PLDs that will feature later in our discussion.
Definition 4.
We denote by $\mathcal{L}_{\mathrm{all}}$ the set of all PLDs and by $\mathcal{L}_{\mathrm{sym}}:=\{\operatorname{PLD}(P\parallel Q):(P,Q)\text{ symmetric}\}$ the set of all symmetric PLDs.
Definition 5.
For $\varepsilon\geq 0$, $\delta\in[0,1]$, let
$R_{\varepsilon,\delta}:=\operatorname{PLD}(\delta 1_{\bot}+(1-\delta)\textrm{Bern}(1/(1+e^{\varepsilon}))\parallel\delta 1_{\top}+(1-\delta)\textrm{Bern}(1/(1+e^{-\varepsilon})))$
We say that $L\in\mathcal{L}_{\mathrm{all}}$ satisfies $(\varepsilon,\delta)$-DP if $L\preceq R_{\varepsilon,\delta}$ or $\varepsilon$-DP if $L\preceq R_{\varepsilon}:=R_{\varepsilon,0}$. We denote $\mathcal{L}_{\mathrm{pure}}:=\{R_{\varepsilon}:\varepsilon\geq 0\}$ and $\mathcal{L}_{\mathrm{approx}}:=\{R_{\varepsilon,\delta}:\varepsilon\geq 0,\delta\in[0,1]\}$.
Definition 6.
For $\mu\geq 0$, we define $G_{\mu}:=\operatorname{PLD}(N(0,1)\parallel N(\mu,1))$ and say that $L\in\mathcal{L}_{\mathrm{all}}$ satisfies $\mu$-GDP when $L\preceq G_{\mu}$. We denote $\mathcal{L}_{\mathrm{gdp}}:=\{G_{\mu}:\mu\geq 0\}$.
Definition 7.
Lastly, we say that $L\in\mathcal{L}_{\mathrm{all}}$ satisfies $(\alpha,\rho$)-RDP if $\mathop{\mathbb{E}}_{Z\sim L}[e^{(\alpha-1)Z}]\leq e^{(\alpha-1)\rho}$.
All three forms of privacy offer simple composition guarantees: For pure DP, $R_{\varepsilon_{1}}\oplus R_{\varepsilon_{2}}\preceq R_{\varepsilon_{1}+\varepsilon_{2}}$ Dwork et al. (2006b), for GDP $G_{\mu_{1}}\oplus G_{\mu_{2}}=G_{\sqrt{\mu_{1}^{2}+\mu_{2}^{2}}}$ Dong et al. (2019), and given $L_{1}$ satisfying $(\alpha,\rho_{1})$-RDP and $L_{2}$ satisfying $(\alpha,\rho_{2})$-RDP, $L_{1}\oplus L_{2}$ satisfies $(\alpha,\rho_{1}+\rho_{2})$-RDP Mironov (2017). More discussion of symmetry can be found in Appendix˜C.
Closely related to the PLD is the hockey-stick curve: a convex reparameterization of the privacy profile, which is in turn a generalization of $(\varepsilon,\delta)$-DP.
@@
For convenience, we write $H_{x}(M):=H_{x}(M(D_{1})\parallel M(D_{2}))$ and sometimes $h_{M}(x):=H_{x}(M)$ for the hockey-stick curve of a mechanism $M$. The hockey-stick curve is very closely related to the classical $(\varepsilon,\delta)$ form of differential privacy: a mechanism $M$ satisfies $(\varepsilon,\delta)$-DP exactly when $H_{e^{\varepsilon}}(M)\leq\delta$. In general, hockey-stick curves are characterized by a few simple conditions.
Proposition 4 (Zhu et al. (2022) Lemma 9).
A curve $h:\mathbb{R}^{\times}\to[0,1]$ is the hockey-stick curve of some $(P,Q)$ iff $h$ is convex and decreasing with $\lim_{x\to 0}h(x)=1$ and $h(x)\geq 1-x$ for all $x\in\mathbb{R}^{\times}$.
Hockey-stick curves are also closely related to PLDs and can be fully recovered from the PLD via the following formula. The proof is in Appendix˜A. Note that we denote $(t)_{+}:=\max\{0,t\}$.